# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v1

@@ 2.3 Sup-irreducible functions and a Krein-Milman theorem
Let $\mathcal{G}$ be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ satisfying 1 and assume that $\mathcal{G}$ is stable by decreasing nets. For all $x\in\mathscr{X}$, define $S(x):=\{u(\cdot)\in\mathcal{G}\,|\,u(x)\geq 0\}$, and let $\operatorname{Min}(S(x))$ be the set of the minimal elements of $S(x)$. Then $\operatorname{Min}(S(x))$ is non-empty and, for any $u\in\operatorname{Min}(S(x))$ and $\lambda\in\mathbb{R}$, $u(\cdot)+\lambda$ is sup-irreducible. Moreover we have that these functions sup-generate $\mathcal{G}$, i.e. $\mathcal{G}=\{\sup_{x\in\mathscr{X},u\in\operatorname{Min}(S(x))}u(\cdot)+a_{x,u}\,|\,a_{x,u}\in\mathbb{R}_{\bot}\}$.
@@
Fix $f\in\mathcal{G}$. Since $\mathcal{G}$ is stable by decreasing nets, for every $x\in\mathscr{X}$ we can select $u\in\operatorname{Min}(S(x))$ such that for all $x^{\prime}\in\mathscr{X}$, $f(x^{\prime})-f(x)\geq u(x^{\prime})$. In other words, for all $x^{\prime}\in\mathscr{X}$,
@@
We say that $b:\mathscr{X}\times\mathscr{X}\to\overline{\mathbb{R}}$ is strictly tropically monotone if $b(x,x)+b(y,y)\geq b(x,y)+b(y,x)$ for all $x,y\in\mathscr{X}$, with equality iff $x=y$.
@@
The set $Z$ is nonempty. It is also inductive since for any decreasing net $(\mathscr{Z}_{\alpha})_{\alpha\in\mathcal{A}}\in Z^{\mathscr{A}}$ we have that $\bigcap_{\alpha\in\mathcal{A}}\mathscr{Z}_{\alpha}\in Z$ and the decreasing sequence of nonempty compact sets is nonempty. Indeed by definition of $\bar{B}^{\circ}$, the function $(\bar{B}^{\circ}f=\sup_{x\in\mathscr{X}}b(x,\cdot)-f(x))$ is the supremum of continuous functions over $\mathscr{Y}$, so it is lower semicontinuous. Hence the function $(b(x,\cdot)-\bar{B}^{\circ}f)$ is upper semicontinuous. By compacity
@@
Fix for each $x$, one such $y_{\alpha,x}$. Let $\mathscr{Z}_{\cap}:=\bigcap_{\alpha\in\mathcal{A}}\mathscr{Z}_{\alpha}$. Since the net is decreasing for the inclusion, $(y_{\alpha,x})_{\alpha\geq\alpha_{0}}\subset\mathscr{Z}_{\alpha_{0}}$, which is compact, so w.l.o.g. $y_{\alpha,x}\to y_{x}\in\mathscr{Z}_{\alpha_{0}}$. As this holds for any $\alpha_{0}$, by considering subnets, we obtain that $y_{x}\in\mathscr{Z}_{\cap}$. As $(b(x,\cdot)-\bar{B}^{\circ}f)$ is upper semicontinuous
@@
Since $f$ is sup-irreducible, it is equal to one of the two functions, but this contradicts the minimality of $\bar{\mathscr{Z}}$. Hence $\bar{\mathscr{Z}}$ contains a single element, and $f=b(\cdot,y)-\bar{B}^{\circ}f(y)$ for $y\in\bar{\mathscr{Z}}$. We showed in 25 that the $u(\cdot)+\lambda$ were sup-irreducible.
Assume now that $\operatorname*{Rg}(B)$ is stable by decreasing nets, $\mathscr{X}=\mathscr{Y}$ and $b$ is strictly tropically monotone. Fix $x\in\mathscr{X}$, and take $u(\cdot)\in\operatorname{Min}(S(x))$ such that $b(\cdot,x)-b(x,x)\geq u(\cdot)$. Since $u\in\operatorname*{Rg}(B)$, for any $\epsilon>0$, we can find $y_{x,\epsilon}$ such that $u(x)\geq b(x,y_{x,\epsilon})-(\bar{B}^{\circ}u)(y_{x,\epsilon})\geq u(x)-\epsilon=-\epsilon$. So
@@
The second part of 26 shows that considering a subset $\bar{\mathscr{Y}}\subset\mathscr{Y}$ such that $\{u(\cdot)+\lambda\}_{x\in X,\lambda\in\overline{\mathbb{R}},u\in\operatorname{Min}(S(x))}=\{b(\cdot,y)+\lambda\}_{y\in\bar{\mathscr{Y}},\lambda\in\overline{\mathbb{R}}}$ would give a sup-irreducible family $\{b(\cdot,y)\}_{y\in\bar{\mathscr{Y}}}$ of $\operatorname*{Rg}(B)$, which sup-generates it owing to 25. However we cannot ensure that $\bar{\mathscr{Y}}$ is compact, a problem that obviously does not happen in finite dimensions. Hence we will make sup-irreducibility an assumption.
We say that a kernel $b:\mathscr{X}\times\mathscr{Y}\to\mathbb{R}$ is fully-reduced if, for all $x\in\mathscr{X}$ and $y\in\mathscr{Y}$, $b(x,\cdot)$ and $b(\cdot,y)$ are sup-irreducible and if, for all $x_{0},x_{1}\in\mathscr{X}$, $y_{0},y_{1}\in\mathscr{Y}$ and $\lambda\in\mathbb{R}$, the implications $(b(\cdot,y_{0})=b(\cdot,y_{1})+\lambda\implies y_{0}=y_{1})$ and $(b(x_{0},\cdot)=b(x_{1},\cdot)+\lambda\implies x_{0}=x_{1})$ both hold, i.e. the rows and columns of $B$ are pairwise independent. It is sometimes possible to show this property directly as in 43 for convex lower semicontinuous functions. Otherwise the second part of 26 provides a testable inequality, satisfied for instance for any $p>0$ by $b=-d(x,y)^{p}$ defined over a metric space $(\mathscr{X},d)$.
@@
In particular, evaluating at $y^{\prime}=y^{\prime}_{0}$, by definition of $b^{\prime}$ and $c^{\prime}$, we have that $\psi^{\prime}(\cdot)=-\varphi^{\prime}(y^{\prime}_{0})$, so $c^{\prime}(x^{\prime},y^{\prime}_{0})=b^{\prime}(\tau(x^{\prime}),\sigma(y^{\prime}_{0}))$. Let $(x^{\prime}_{k})_{k\in\mathbb{N}}$ be a sequence converging to some $x^{\prime}\in\mathscr{X}^{\prime}$. Assume that $(\tau(x^{\prime}_{k}))_{k\in\mathbb{N}}$ has two subsequences converging respectively to $x_{0}$ and $x_{1}\in\mathscr{X}$. Then taking the limit, since the kernels are continuous we obtain that
@@
However since the kernel $b$ is fully-reduced and $\sigma$ is bijective, we must have $x_{0}=x_{1}$. Consequently $\tau$ is continuous. Symmetrizing the proof, so is $\tau^{-1}$ and we get that $\tau$ and $\sigma$ are homeomorphisms. As we already established (23), since $b$ and $c$ are continuous and the bijections $\sigma$ and $\tau$ are continuous, so are $\psi$ and $\varphi$.
@@
Since by 9, the kernel of the operator $\bar{C}$ is unique, we have that
@@
$\sup_{y^{\prime}\in\mathscr{Y}^{\prime}}\left[-\sup_{z\in\mathscr{X}}[\varphi(y^{\prime})+b(z,\sigma(y^{\prime}))-g(z)]+\inf_{x\in\mathscr{X}}[f(x)+\varphi(y^{\prime})+b(x,\sigma(y^{\prime}))]\right]\\ =\sup_{y\in\mathscr{X}}-\bar{B}g(y)+\inf_{x\in\mathscr{X}}f(x)+b(x,y).$
@@ 3 Studying (max,+) and order isomorphisms over specific sets
Applying the results of Section 2, and in particular 21, we now characterize the (max,+)-isomorphisms for some given sets of functions: lower semicontinuous or continuous, 1-Lipschitz or Lipschitz with a (weak) locally compact or complete metric, lower semicontinuous and convex on a Hausdorff vector space. For some of these sets we can even characterize the order isomorphisms in great generality.
@@ 3.1 Indicator functions and lower semicontinuous functions
The Banach-Stone theorem is a classical result of linear functional analysis. In this section, we give a (max,+) version of it. The following characterization (25) appeared in [KM97, Theorem 2.2] for (max,+)-isomorphisms over the set of continuous functions on a complete locally compact set $\mathscr{X}$. Inf-irreducible points were not used in the proof of Maslov and Kolokoltsov, the latter relied instead on a representation theorem of integral operators which we circumvented. We also dropped their requirements on $\mathscr{X}$.
@@
where $g:\mathscr{X}\rightarrow\mathbb{R}$ is a continuous function and $\phi:\mathscr{X}\rightarrow\mathscr{Y}$ is a homeomorphism. The same holds if lower semicontinuous is replaced by continuous, or if the functions are restricted to be proper.
@@
Concerning the set of continuous functions $\mathcal{C}(\mathscr{X},\mathbb{R})$, by 15, since its sup-closure is the set of proper lower semicontinuous functions, the (max,+)-isomorphisms between spaces of continuous functions are of the form (25). The converse part is immediate.∎
The following result for pointwise dense subsets of lower semicontinuous functions encompasses and greatly simplifies [LT16, Theorem 4.5]. The latter is written for some specific vector subspaces of continuous functions, named “near vector lattices” by the authors, and encompassing (locally) Lipschitz or uniformly continuous (bounded) functions. Near vector lattices $\mathcal{G}$ require that one can separate closed sets from points with functions of $\mathcal{G}$, such $\mathcal{G}$ are thus pointwise dense in continuous functions and are covered by our theorem. [^2] A short proof for this fact is that, if for all $x\in\mathscr{X}$ and all closed $F\subset\mathscr{X}$ not containing $x$, there exists $g\in\mathcal{G}$ such that $g(x)\neq 0$ and $g_{|F}=0$, then, for all finite $F\subset\mathscr{X}$ and all $x\in F$, there exists a function $g_{F,x}\in\mathcal{G}$ such that its restriction to $F$ is $1$ if $y=x$ and 00 if $y\in F\backslash\{x\}$. However $\operatorname*{span}\{g_{F,x}\,|\,F\subset\mathscr{X}\text{ finite, }x\in F\}$ is pointwise dense in $C(\mathscr{X},\mathbb{R})$, so $\mathcal{G}$ is also pointwise dense.
@@
where $\phi$ is a homeomorphism and $g:\mathscr{X}\times\overline{\mathbb{R}}\rightarrow\overline{\mathbb{R}}$ is jointly lower semicontinuous, and $g(x,\cdot)$ bijective and increasing for all $x\in\mathscr{X}$ with inverse $g^{1}(x,\cdot)$ such that $g^{1}(\cdot,\cdot)$ is also jointly lower semicontinuous. The same holds if lower semicontinuous is replaced by continuous.
More generally, if $\mathcal{G}$ (resp. $\mathscr{F}$) is a pointwise-dense subset of the lower semicontinuous functions over a Hausdorff topological space $\mathscr{X}$ (resp. $\mathscr{Y}$) then (26) holds necessarily, with $\phi$ and $g$ as above.
@@
The $\mathcal{G}$-relatively-inf-irreducible points of lower semicontinuous functions are the Dirac masses $(\delta_{x}+\lambda)_{x\in\mathscr{X},\lambda\in\overline{\mathbb{R}}}$. Indeed if $g,h\in\overline{\mathbb{R}}^{\mathscr{Y}}$ are such that $\delta_{x}=\inf\nolimits^{\mathcal{G}}(g,h)$, then $\delta_{x}\leq g$ and $\delta_{x}\leq h$, so both are Dirac masses $\delta_{x}+\lambda$, one being equal to $\delta_{x}$. Conversely take $\mathcal{G}$-relatively-inf-irreducible $f\in\mathcal{G}$. It is necessarily inf-irreducible by 5. Assume there exists $x_{0}\neq x_{1}$ such that $f(x_{0})<\infty$ and $f(x_{1})<\infty$, then fix a closed subset $F\subset\mathscr{X}$ such that $x_{0}\in F$, $x_{0}\notin\overline{\mathscr{X}\backslash F}$ and $x_{1}\notin F$. Hence $f=\inf(f+\delta_{F},f+\delta_{\overline{\mathscr{X}\backslash F}})$, both of these functions being different than $f$. Thus $f=\delta_{x}+\lambda$ for some $\lambda\in\overline{\mathbb{R}}$.
@@
Composing with $g_{y}$, we obtain $Jf(\phi(y))=g_{y}(f(y))$. Renaming variables to write $Jf(x)=g_{x}(f(\phi(x)))$ and $J^{-1}f(x)=g^{1}_{x}(f(\phi^{-1}(x)))$, we get the desired result (26) with now $g_{x}(g^{1}_{\phi(x)}(f(x)))=f(x)$ after renaming. Moreover $J(y\mapsto\lambda)(x)=g(x,\lambda)$ is lower semicontinuous. The same holds for $g^{1}$. Thus $g(\cdot,\cdot)$ is separately lower semicontinuous with $g(x,\cdot)$ increasing. Consequently, by [Gru19, Theorem 1], $g(\cdot,\cdot)$ is thus jointly lower semicontinuous.333 [Gru19, Theorem 1] is written for continuity rather than lower semicontinuity, but taking $T=T_{1}=\mathbb{R}$ in Grushka’s notations, and setting $\tau_{2}=+\infty$ in his proof, the result straightforwardly extends to lower semicontinuous functions since they have almost the same characterization through neighborhoods as continuous functions. By similar arguments, so is $g^{1}$. Using only the indicator functions, we thus obtained the intermediary result that
@@
with $\phi$ bijective and $g$ jointly lower semi-continuous, with $g$ continuous and increasing in the second variable.
@@
Concerning the set of continuous functions $\mathcal{C}(\mathscr{X},\mathbb{R})$, by 15, since its sup-closure is the set of proper lower semicontinuous functions, the order isomorphisms between spaces of continuous functions are of the form (27). In all the proof above, we just replace lower semicontinuous by continuous with no other change to derive the result. Similarly the last statement is a direct consequence of 15. ∎
If $\mathscr{X}$ is discrete, then $-\operatorname*{Id}$ is an anti-involution over (non proper) lower semicontinuous functions. However if the space has no isolated points then there are no anti-involutions.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be the space of lower semicontinuous functions over a set $\mathscr{X}$ (resp. $\mathscr{Y}$). If there exists an order anti-isomorphism $T$ from $\mathscr{F}$ to $\mathcal{G}$, then $\mathscr{X}$ (resp. $\mathscr{Y}$) is in bijection with the set $I_{\mathscr{X}}$ of the isolated points of $\mathscr{Y}$ (resp. $\mathscr{X}$) and there exists a bijection $\phi:\mathscr{X}\to\mathscr{Y}$. In particular, if $\mathscr{X}$ or $\mathscr{Y}$ do not have isolated points, or if $I_{\mathscr{X}}$ and $I_{\mathscr{Y}}$ do not have the same cardinality, then there is no order anti-isomorphism.
The intuition here is that anti-involutions send relative-inf-irreducible points onto sup-irreducible points by 13. However the sup-irreducible points of lower semicontinuous functions are only the $\delta^{\bot}_{x}$ of isolated $x$.

### v2

@@ 2.3 Sup-irreducible functions and a Krein–Milman theorem
Let $\mathcal{G}$ be a complete subspace of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ and assume that $\mathcal{G}$ is stable by decreasing nets. For all $x\in\mathscr{X}$, define $S(x):=\{u(\cdot)\in\mathcal{G}\,|\,u(x)\geq 0\}$, and let $\operatorname{Min}(S(x))$ be the set of the minimal elements of $S(x)$. Then $\operatorname{Min}(S(x))$ is non-empty and, for any $u\in\operatorname{Min}(S(x))$ and $\lambda\in\mathbb{R}$, $u(\cdot)+\lambda$ is sup-irreducible. Moreover we have that these functions sup-generate $\mathcal{G}$, i.e. $\mathcal{G}=\{\sup_{x\in\mathscr{X},u\in\operatorname{Min}(S(x))}u(\cdot)+a_{x,u}\,|\,a_{x,u}\in\mathbb{R}_{\bot}\}$.
@@
Fix $f\in\mathcal{G}$. Since $\mathcal{G}$ is stable by decreasing nets, for every $x\in\mathscr{X}$ we can select $u\in\operatorname{Min}(S(x))$ such that for all $x^{\prime}\in\mathscr{X}$, $f(x^{\prime})-f(x)\geq u(x^{\prime})$color=blue!10,size=] PCA: vérifier cette phrase. In other words, for all $x^{\prime}\in\mathscr{X}$,
@@
We say that $b:\mathscr{X}\times\mathscr{X}\to\overline{\mathbb{R}}$ is strictly tropically monotone if $b(x,x)+b(y,y)\geq b(x,y)+b(y,x)$ for all $x,y\in\mathscr{X}$, with equality iff $x=y$. This holds for $(\mathscr{X},d)$ a metric space with $b=-d$ or for $X$ a Hilbert space with $b=\langle\cdot,\cdot\rangle$.
@@
The set $Z$ is nonempty since $\mathscr{Y}\in Z$. It is also inductive since for any decreasing net $(\mathscr{Z}_{\alpha})_{\alpha\in\mathcal{A}}\in Z^{\mathscr{A}}$ we have that $\bigcap_{\alpha\in\mathcal{A}}\mathscr{Z}_{\alpha}\in Z$ and the decreasing sequence of nonempty compact sets is nonempty. Indeed by definition of $\bar{B}^{\circ}$, the function $(\bar{B}^{\circ}f=\sup_{x\in\mathscr{X}}b(x,\cdot)-f(x))$ is the supremum of continuous functions over $\mathscr{Y}$, so it is lower semicontinuous. Hence the function $(b(x,\cdot)-\bar{B}^{\circ}f)$ is upper semicontinuous. By compactness
@@
Fix for each $x$, one such $y_{\alpha,x}$. Let $\mathscr{Z}_{\cap}:=\bigcap_{\alpha\in\mathcal{A}}\mathscr{Z}_{\alpha}$. Since the net is decreasing for the inclusion, $\{y_{\alpha,x}\}_{\alpha\geq\alpha_{0}}\subset\mathscr{Z}_{\alpha_{0}}$, which is compact, so w.l.o.g. $y_{\alpha,x}\to y_{x}\in\mathscr{Z}_{\alpha_{0}}$. As this holds for any $\alpha_{0}$, by considering subnets, we obtain that $y_{x}\in\mathscr{Z}_{\cap}$. As $(b(x,\cdot)-\bar{B}^{\circ}f)$ is upper semicontinuouscolor=red!30]@SG: j’ai besoin que tu vérifies les preuves de la page 15. Très dangereux avec les subnets
@@
Since $f$ is sup-irreducible, it is equal to one of the two functions, but this contradicts the minimality of $\bar{\mathscr{Z}}$. Hence $\bar{\mathscr{Z}}$ contains a single element, and $f=b(\cdot,y)-\bar{B}^{\circ}f(y)$ for $y\in\bar{\mathscr{Z}}$.color=blue!10,size=] PCA: je ne pense qu’on ait besoin de "stable by decreasing nets" jusqu’ici
Assume now that $\operatorname*{Rg}(B)$ is stable by decreasing nets, $\mathscr{X}=\mathscr{Y}$ and $b$ is strictly tropically monotone. We showed in 25 that the functions $u(\cdot)+\lambda$ were sup-irreducible, which gives one inclusion. To show the reverse inclusion, fix $x\in\mathscr{X}$, and take $u(\cdot)\in\operatorname{Min}(S(x))$ such that $b(\cdot,x)-b(x,x)\geq u(\cdot)$. Since $u\in\operatorname*{Rg}(B)$, for any $\epsilon>0$, we can find $y_{x,\epsilon}$ such that $u(x)\geq b(x,y_{x,\epsilon})-(\bar{B}^{\circ}u)(y_{x,\epsilon})\geq u(x)-\epsilon=-\epsilon$. So
@@
When $\mathcal{G}$ is stable by decreasing nets, the first part of 26 shows that considering a subset $\bar{\mathscr{Y}}\subset\mathscr{Y}$ such that $\{u(\cdot)+\lambda\}_{x\in X,\lambda\in\overline{\mathbb{R}},u\in\operatorname{Min}(S(x))}=\{b(\cdot,y)+\lambda\}_{y\in\bar{\mathscr{Y}},\lambda\in\overline{\mathbb{R}}}$ would give a sup-irreducible family $\{b(\cdot,y)\}_{y\in\bar{\mathscr{Y}}}$ of $\operatorname*{Rg}(B)$, which sup-generates it owing to 25. However we cannot ensure that $\bar{\mathscr{Y}}$ is compact, a problem that obviously does not happen in finite dimensions. Hence we will make sup-irreducibility an assumption.
We say that a kernel $b:\mathscr{X}\times\mathscr{Y}\to\mathbb{R}$ is fully-reduced if, for all $x\in\mathscr{X}$ and $y\in\mathscr{Y}$, $b(x,\cdot)$ and $b(\cdot,y)$ are sup-irreducible and if, for all $x_{0},x_{1}\in\mathscr{X}$, $y_{0},y_{1}\in\mathscr{Y}$ and $\lambda\in\mathbb{R}$, the implications $(b(\cdot,y_{0})=b(\cdot,y_{1})+\lambda\implies y_{0}=y_{1})$ and $(b(x_{0},\cdot)=b(x_{1},\cdot)+\lambda\implies x_{0}=x_{1})$ both hold, i.e. the rows and columns of $B$ are pairwise independent. It is sometimes possible to show this property directly as in 41 for convex lower semicontinuous functions. Otherwise the second part of 26 provides a testable inequality, satisfied for instance for any $p>0$ by $b=-d(x,y)^{p}$ over a metric space $(\mathscr{X},d)$. For $p\in(0,1]$, $b$ defines the inf-stable $\operatorname*{Rg}(B)$ of Hölder functions, which are indeed stable by decreasing nets. If $\mathscr{Y}$ is finite, $\operatorname*{Rg}(B)$ is also stable by decreasing nets.color=blue!10,size=] PCA: vérifier la dernière phrase
@@
In particular, evaluating at $y^{\prime}=y^{\prime}_{0}$, by definition of $b^{\prime}$ and $c^{\prime}$, we have that $\psi^{\prime}(\cdot)=-\varphi^{\prime}(y^{\prime}_{0})$, so $c^{\prime}(\cdot,y^{\prime}_{0})=b^{\prime}(\tau(\cdot),\sigma(y^{\prime}_{0}))$. Let $(x^{\prime}_{\alpha})_{\alpha\in\mathscr{A}}$ be a net converging to some $x^{\prime}\in\mathscr{X}^{\prime}$. Assume that $(\tau(x^{\prime}_{\alpha}))_{\alpha\in\mathscr{A}}$, which belongs to the compact set $\mathscr{X}$, has two subnets converging respectively to $x_{0}$ and $x_{1}\in\mathscr{X}$. Then taking the limit, since the kernels are continuous we obtain that
@@
However since the kernel $b$ is fully-reduced and $\sigma$ is bijective, we must have $x_{0}=x_{1}$. Consequently, all converging subnets have the same limit $x_{0}$. Similarly to [AB06, Lemma 2.17], if $(\tau(x^{\prime}_{\alpha}))_{\alpha\in\mathscr{A}}$ did not converge to $x_{0}$, then it would have a subnet staying away from a neighborhood of $x_{0}$, but by compactness this subnet has itself a converging subnet to $x_{0}$.color=blue!10,size=] PCA: @SG: vérifie stp We have thus shown that $\tau$ is continuous. Symmetrizing the proof, so is $\tau^{-1}$ and we get that $\tau$ and $\sigma$ are homeomorphisms. As we already established (25), since $b$ and $c$ are continuous and the bijections $\sigma$ and $\tau$ are continuous, so are $\psi$ and $\varphi$.
@@
Since by 9, the kernel of the operator $\bar{C}^{\circ}$ is unique, we have that
@@
$\sup_{y^{\prime}\in\mathscr{Y}^{\prime}}\left[-\sup_{z\in\mathscr{X}}[\varphi(y^{\prime})+b(z,\sigma(y^{\prime}))-g(z)]+\inf_{x\in\mathscr{X}}[f(x)+\varphi(y^{\prime})+b(x,\sigma(y^{\prime}))]\right]\\ =\sup_{y\in\mathscr{Y}}-\bar{B}g(y)+\inf_{x\in\mathscr{X}}f(x)+b(x,y).$
@@ 3 Studying (max,+) and order isomorphisms over specific sets
Applying the results of Section 2, and in particular 21, we now characterize the (max,+)-isomorphisms for some given sets of functions: i) lower semicontinuous or continuous, ii) 1-Lipschitz with a (weak) locally compact or complete metric, iii) lower semicontinuous and convex. For some of these sets we can even characterize the order isomorphisms in great generality.
@@ 3.1 Indicator functions and lower semicontinuous functions
The Banach-Stone theorem is a classical result of linear functional analysis. In this section, we give a (max,+) version of it. The following characterization (27) appeared in [KM97, Theorem 2.2] for (max,+)-isomorphisms over sets of continuous functions on separable locally compact topological spaces with the functions diverging to $\infty$ outside of a compact set, see [KM97, Section 1.4, p34] for their notations. Inf-irreducible points were not used in the proof of Maslov and Kolokoltsov, the latter relied instead on a representation theorem of integral operators which we circumvented.
@@
where $g:\mathscr{X}\rightarrow\mathbb{R}$ is a continuous function and $\phi:\mathscr{X}\rightarrow\mathscr{Y}$ is a homeomorphism. For $\mathscr{X}$ and $\mathscr{Y}$ compact, the same holds if lower semicontinuous is replaced by continuous, or if the functions are restricted to be proper.
@@
Concerning the set of continuous functions $\mathcal{C}(\mathscr{X},\mathbb{R})$, for $\mathscr{X}$ compact, by 15, since its sup-closure is the set of proper lower semicontinuous functions, the (max,+)-isomorphisms between spaces of continuous functions are of the form (27). The converse part is immediate.∎
@@
where $\phi$ is a homeomorphism and $g:\mathscr{X}\times\overline{\mathbb{R}}\rightarrow\overline{\mathbb{R}}$ is jointly lower semicontinuous, and $g(x,\cdot)$ bijective and increasing for all $x\in\mathscr{X}$ with inverse $g^{1}(x,\cdot)$ such that $g^{1}(\cdot,\cdot)$ is also jointly lower semicontinuous. For $\mathscr{X}$ and $\mathscr{Y}$ compact, the same holds if lower semicontinuous is replaced by continuous and $J$ satisfies (13).
Concerning lattice isomorphisms $J$ of continuous functions over compact Hausdorff spaces, assuming $J$ to be continuous under pointwise convergence, which is weaker than (13), formula (28) was first shown in [Kap48, Theorem 1] with an elegant proof based on characterizing “prime ideals” of $C(\mathscr{X},\mathbb{R})$. When the continuity assumption is dropped, formula (28) still holds on a $G_{\delta}$-dense set [CS08, Theorem 1, Lemma 4], which encompasses the results of Kaplansky. In [LT16, Theorem 4.5], the same expression was obtained for isomorphims of some specific vector subspaces of continuous functions on a metric space, named “near vector lattices” by the authors, and encompassing (locally) Lipschitz or uniformly continuous (bounded) functions. Compactifications of $\mathscr{X}$ and continuity of $J$, which is not automatic, are key elements in this approach. They do not appear when focusing on lower semicontinuous functions.
@@
The $\mathcal{G}$-relatively-inf-irreducible points of lower semicontinuous functions are the Dirac masses $(\delta_{x}+\lambda)_{x\in\mathscr{X},\lambda\in\overline{\mathbb{R}}}$. Indeed if $g,h\in\overline{\mathbb{R}}^{\mathscr{Y}}$ are such that $\delta_{x}=\inf\nolimits^{\mathcal{G}}(g,h)$, then $\delta_{x}\leq g$ and $\delta_{x}\leq h$, so both are Dirac masses $\delta_{x}+\lambda$, one being equal to $\delta_{x}$. Conversely take a $\mathcal{G}$-relatively-inf-irreducible $f\in\mathcal{G}$. It is necessarily inf-irreducible by 5. Assume there exists $x_{0}\neq x_{1}$ such that $f(x_{0})<\infty$ and $f(x_{1})<\infty$, then fix a closed subset $F\subset\mathscr{X}$ such that $x_{0}\in F$, $x_{0}\notin\overline{\mathscr{X}\backslash F}$ and $x_{1}\notin F$. Hence $f=\inf(f+\delta_{F},f+\delta_{\overline{\mathscr{X}\backslash F}})$, both of these functions being different than $f$. Thus $f=\delta_{x}+\lambda$ for some $\lambda\in\overline{\mathbb{R}}$.
@@
Composing with $g_{y}$, we obtain $Jf(\phi(y))=g_{y}(f(y))$. Renaming variables to write $Jf(x)=g_{x}(f(\phi(x)))$ and $J^{-1}f(x)=g^{1}_{x}(f(\phi^{-1}(x)))$, we get the desired result (28) with now $g_{x}(g^{1}_{\phi(x)}(f(x)))=f(x)$ after renaming. Moreover $J(y\mapsto\lambda)=g(\cdot,\lambda)$ is lower semicontinuous. The same holds for $g^{1}$. Thus $g(\cdot,\cdot)$ is separately lower semicontinuous with $g(x,\cdot)$ increasing. Consequently, by [Gru19, Theorem 1], $g(\cdot,\cdot)$ is jointly lower semicontinuous.444 [Gru19, Theorem 1] is written for continuity rather than lower semicontinuity, but taking $T=T_{1}=\mathbb{R}$ in Grushka’s notations, and setting $\tau_{2}=+\infty$ in his proof, the result straightforwardly extends to lower semicontinuous functions since they have almost the same characterization through neighborhoods as continuous functions. By similar arguments, so is $g^{1}$. Using only the indicator functions, we have obtained the intermediary result that
@@
with $\phi$ bijective and $g$ jointly lower semicontinuous, with $g$ continuous and increasing in the second variable.
@@
Concerning the set of continuous functions $\mathcal{C}(\mathscr{X},\mathbb{R})$, for $\mathscr{X}$ compact and $J$ satisfying (13), by 15, since its sup-closure is the set of proper lower semicontinuous functions, the order isomorphisms between spaces of continuous functions are of the form (29). In all the proof above, we just replace lower semicontinuous by continuous with no other change to derive the result applying 15.∎
About anti-involutions, if $\mathscr{X}$ is discrete, then $-\operatorname*{Id}$ is an anti-involution over (non proper) lower semicontinuous functions. However if the space has no isolated points then there are no anti-involutions.
@@
Let $\mathcal{G}$ (resp. $\mathscr{F}$) be the space of lower semicontinuous functions over a set $\mathscr{X}$ (resp. $\mathscr{Y}$). If there exists an order anti-isomorphism $T$ from $\mathscr{F}$ to $\mathcal{G}$, then $\mathscr{X}$, the set $I_{\mathscr{X}}$ of the isolated points of $\mathscr{X}$, $\mathscr{Y}$ and $I_{\mathscr{Y}}$ are all in bijection. In particular, if $\mathscr{X}$ or $\mathscr{Y}$ do not have isolated points, or if $I_{\mathscr{X}}$ and $I_{\mathscr{Y}}$ do not have the same cardinality, then there is no order anti-isomorphism.
The intuition here is that anti-isomorphisms send relative-inf-irreducible points onto sup-irreducible points by 13. However the sup-irreducible points of lower semicontinuous functions are only the $\delta^{\bot}_{x}$ of isolated $x$.